# Beyond Local Sharpness: Communication-Efficient Global Sharpness-aware Minimization

### v1

@@ Appendix A Algorithm
Algorithm 1 summarizes FedGloSS, using as an example Sgd or Sam as local optimizers, differently highlighted. The comment colors replicate those of Fig. 3.
@@
22:     $\sigma^{t+1}=\sigma^{t}-\frac{1}{\beta|\mathcal{C}|}\sum_{k\in\mathcal{C}}(\operatorname{\boldsymbol{w}}_{k}^{t}-\operatorname{\boldsymbol{w}}^{t})$triangleright Update global dual variable
@@
Appendix B The Benefits of FedGloSS
@@ B.1 Achieving Consistency on Local and Global Sharpness with FedGloSS
This section completes the analyses presented in Sections 4 and 6.2, showing how FedGloSS guides towards consistent local and global flat loss landscapes. Fig. 10 extends Figs. 5(a) and 5(b) from the main paper with the 3 remaining clients (out of 5) selected in the last round, comparing the behavior of clients’ models trained with FedGloSS from the local and global perspectives. These results highlight the effectiveness of FedGloSS in achieving local models with consistent local-global behavior. Indeed, these models end up in flat regions both in local and global loss landscapes. The comparison with FedSam (net surface) further demonstrates the effectiveness of using our global flatness-aware approach.
@@
Fig. 11 instead extends Figs. 5(c) and 5(d), offering a comparative analysis of FedSmoo’s clients’ models w.r.t. FedGloSS (net landscape). The local solutions found by FedGloSS achieve flatter, better (i.e., lower loss) and more consistent convergence points in the global loss landscape w.r.t. our main competitor FedSmoo.
@@ B.2 Achieving Flatter Global Minima
Table 8 reports the values of the maximum Hessian eigenvalues, as already shown in Fig. 7 in the main paper. First, as expected, we note that Sam-based methods achieve flatter minima w.r.t. the counterpart. Notably, our main competitor FedSmoo presents higher sharpness than FedSam in the simpler Cifar10, regardless of the data distribution. In addition, FedGloSS with local Sam achieves the lowest sharpness (i.e., lowest $\lambda_{1}$) on all configurations, outperforming the state of the art and, specifically, all sharpness-aware methods.
@@
B.3 Increasing Convergence Speed
Figs. 15 and 16 show the accuracy trends of FedGloSS compared to state-of-the-art methods for heterogeneous FL on Cifar10 and Cifar100 respectively. For a clearer understanding, we distinguish between Sam-based and Sgd-based methods depending on the used local optimizer. For a fair comparison, we report FedSmoo with and without the scheduling of $\rho$ (+wp in the figure). For additional details on the scheduling, refer to Appendix C. FedGloSS consistently achieves the best performances and convergence speedup in each group. In addition, we remind that FedGloSS communicates half the number of bits at each round w.r.t. FedSmoo. Fig. 17 reports the results obtained with ResNet18 on both datasets: FedGloSS consistently achieves the best speed of convergence and final accuracy, with both Sam and Sgd as ClientOpt.
@@ B.4 Effectiveness in Multiple Scenarios and with Several Model Architectures
As already shown in Section 6, FedGloSS outperforms the state of the art in terms of speed of convergence, final performance in accuracy, flatness of the solution and communication efficiency. This remains true across multiple datasets (Cifar100, Cifar10, Landmarks-User-160k) and model architectures (CNN, ResNet18, MobileNetv2). The main paper focuses on results in more challenging heterogeneous FL scenarios. The next paragraph discusses the behavior of FedGloSS in homogeneous settings.
@@
Table 9 evaluates FedGloSS against the main methods FedAvg, FedSam, FedDyn and FedSmoo in homogeneous FL settings. Here, client gradients are naturally more aligned due to reduced client drift (Karimireddy et al., 2020b). We thus test FedGloSS with and without ADMM for global consistency. As expected, FedGloSS achieves similar accuracy with or without ADMM, particularly when using SAM as the local optimizer. However, ADMM facilitates convergence to flatter minima (evidenced by lower $\lambda_{1}$ values) by aligning local and global convergence points. Notably, FedGloSS achieves the flattest minima (lowest $\lambda_{1}$) across both datasets, and the best accuracy on the more complex Cifar100. While FedSmoo achieves slightly higher accuracy on Cifar10, FedGloSS finds a flatter minimum and achieves competitive accuracy with significantly lower communication costs (halved).
@@ B.5 Reducing Communication Costs
Table 10 extends the analysis presented in Section 6.3.4 by evaluating the communication costs across all scenarios considered in this work. Notably, since FedGloSS maintains the per-round communication cost of FedAvg, it remains advantageous even when matching the convergence speed of the best-performing algorithm (FedSmoo), due to its halved communication costs.
@@
Fig. 18 deepens the comparison between FedGloSS and its baseline NaiveFedGloSS, discussed in Section 6.4. In particular, this plot reports the accuracy trends of the two methods, showing that NaiveFedGloSS is slightly faster (1.1$\approx$ 1.1$\times$ 1.1 ×) than FedGloSS in Cifar100, while FedGloSS surpasses the speed of the baseline after $\approx 25\%$ of training in Cifar10. However, both methods reach the same accuracy at the of training. In addition, it is to be noted that FedGloSS halves the communication cost w.r.t. NaiveFedGloSS by transmitting half the number of bits at each round. Additionally, it also reduces the communication cost by half, as it eliminates the need to invoke the clients twice to compute the updates. These insights further support our choice of the efficient strategy of FedGloSS over NaiveFedGloSS.
@@ Appendix C Implementation Details
This section delves into a comprehensive description of the datasets and models utilized throughout this paper, specifying the Deep Learning framework and the hardware employed (Section C.1). Additionally, we present the area of the hyper-parameters’ space explored during the fine-tuning process in order to yield optimal results (Section C.2).
@@ C.1 Datasets and Models
Table 11 provides a comprehensive overview of each dataset’s general statistics. This includes the number of training clients participating in the process and the total number of samples used to construct both the training and test sets.
@@ C.1.1 Cifar10 and Cifar100
We adapted these two well-known image classification datasets to the FL scenario by replicating the splits among clients proposed by (Hsu et al., 2019). Both datasets are split evenly among 100 clients, thus each of them has access to 500 data samples. This partitioning is performed according to a Latent Dirichlet Allocation (LDA) on the labels. In practice, each local dataset follows a multinomial distribution drawn according to a symmetric Dirichlet distribution with concentration parameter $\alpha$. The higher the value of this parameter is, the more the local datasets resemble a homogeneous scenario, in the limit case $\alpha=0$ each client has access to one only class of images. In our experiments we tested $\alpha\in\{0,0.05,100\}$ and $\alpha\in\{0,0.5,1000\}$ for Cifar10 and Cifar100, respectively. Figs. 20(a) and 20(b) show how data is distributed across clients in all the experimental settings for these two datasets. Both datasets are pre-processed by applying random crops and random horizontal flips.
@@
We trained a Convolutional Neural Network (CNN) inspired by the LeNet-5 architecture (LeCun et al., 1998), as proposed by (Hsu et al., 2020). The network comprises two 64-channels convolutional layers, both using $5\times 5$ kernels and followed by $2\times 2$ max-pooling layers. This is succeeded by two fully-connected layers with 384 and 192 units, respectively. The final output layer is adapted to the specific number of classes in the dataset.
To explore deeper and more expressive architectures, we also trained a ResNet18 (He et al., 2015) on Cifar100 with $\alpha=0.5$ and Cifar10 with $\alpha=0.05$. We replaced the standard Batch Normalization layers (Ioffe & Szegedy, 2015) with Group Normalization layers (Wu & He, 2018) due to their demonstrated effectiveness in handling skewed data distributions in FL (Hsieh et al., 2020). We carried out the experiments with the CNN model using PyTorch (Paszke et al., 2019) and the ResNet18 ones using FedJAX (Ro et al., 2021).
@@ C.1.2 Landmarks-User-160k
To achieve a comprehensive understanding of the efficacy of the proposed method, a thorough analysis was undertaken utilizing large-scale real-world datasets. Specifically, we use the Landmarks-User-160k dataset (Hsu et al., 2020), a repository encompassing 164,172 images depicting 2028 distinct landmarks, distributed among 1262 clients.
@@
The model employed for training is a MobileNetV2 (Sandler et al., 2018; Hsu et al., 2020), replacing batch normalization with group normalization layers and pre-trained on ImageNet (Deng et al., 2009). The tests on this dataset were carried out on a cluster of NVIDIA A100 40GB, using our FedJAX codebase.
@@ C.2 Hyperparameters
In Table 12 we report the training hyperparameters associated to each dataset and model pairing. While Table 13 summarizes the hyper-parameters search grid tested for each method (in bold the chosen ones). All runs are averaged over 3 seeds. In addition, following previous works (Acar et al., 2021; Caldarola et al., 2022; 2023), we report the averaged accuracy of the last 100 rounds to reduce the noise typical of heterogeneous FL settings.
@@ Appendix D Flatness Analysis
D.1 Visualizing 1D Loss Landscapes
Fig. 9 in the main paper shows the interpolation of FedGloSS and NaiveFedGloSS’s models. Given their respective weights $\operatorname{\boldsymbol{w}}_{\textsc{FedGloSS}}$ and $\operatorname{\boldsymbol{w}}_{\textsc{NaiveFedGloSS}}$, the interpolation is computed using a coefficient $\gamma$ as
@@ D.1 Visualizing 1D Loss Landscapes
D.2 Visualizing 3D Loss Landscapes
We leverage techniques from (Li et al., 2018) to visualize the loss landscapes of our models. We adapt their code to work with our specific datasets and network architectures. The process involves calculating the loss function along random directions in the parameter space. This is achieved by perturbing the model’s parameters within a defined range. In our visualizations, we constrain these perturbations to occur within the range of $[-1,1]$ for both the $x$ and $y$ axes. To ensure consistent comparisons across models (e.g., as seen in Fig. 6), we utilize the same set of random directions for all models.
@@ D.3 Hessian Eigenvalues for Flatness Measure
Following prior work (Foret et al., 2021; Garipov et al., 2018; Li et al., 2018; Caldarola et al., 2022), we use the spectrum of the Hessian matrix to quantify the flatness of the loss landscape. Here, lower maximum eigenvalues correspond to flatter landscapes, implying less sharpness. To compute these eigenvalues (denoted by $\lambda_{1}$ in the main paper), we employ the stochastic power iteration method (Xu et al., 2018) with a maximum of 20 iterations, referring to the code of Golmant et al. (2018).

### v2

@@ Appendix A Algorithm
Alg. 1 summarizes FedGloSS, using as an example Sgd or Sam as local optimizers, differently highlighted. The comment colors replicate those of Fig. 3.
@@
22:    $\sigma^{t+1}=\sigma^{t}-\frac{1}{\beta|\mathcal{C}|}\sum_{k\in\mathcal{C}}(\operatorname{\boldsymbol{w}}_{k}^{t}-\tilde{\operatorname{\boldsymbol{w}}}^{t})$triangleright Update global dual variable
@@
Appendix B The Benefits on FedGloSS
@@ B.1 Achieving Consistency on Local and Global Sharpness with FedGloSS
This section completes the analyses presented in Secs. 4 and 6.2, showing how FedGloSS guides towards consistent local and global flat loss landscapes. Fig. 10 extends Figs. 4(a) and 4(b) from the main paper with the 3 remaining clients (out of 5) selected in the last round, comparing the behavior of clients’ models trained with FedGloSS from the local and global perspectives. These results highlight the effectiveness of FedGloSS in achieving local models with consistent local-global behavior. Indeed, these models end up in flat regions both in local and global loss landscapes. The comparison with FedSam (net surface) further demonstrates the effectiveness of using our global flatness-aware approach.
Fig. 11 instead extends Figs. 4(c) and 4(d), offering a comparative analysis of FedSmoo’s clients’ models w.r.t. FedGloSS (net landscape). The local solutions found by FedGloSS achieve flatter, better (i.e., lower loss) and more consistent convergence points in the global loss landscape w.r.t. our main competitor FedSmoo.
@@ B.2 Achieving Flatter Global Minima
Tab. 9 reports the values of the maximum Hessian eigenvalues, as already shown in Fig. 7 in the main paper. First, as expected, we note that Sam-based methods achieve flatter minima w.r.t. the counterpart. Notably, our main competitor FedSmoo presents higher sharpness than FedSam in the simpler Cifar10, regardless of the data distribution. In addition, FedGloSS with local Sam achieves the lowest sharpness (i.e., lowest $\lambda_{1}$) on all configurations, outperforming the state of the art and, specifically, all sharpness-aware methods.
@@
B.3 Increasing Convergence Speed
Figs. 15 and 16 show the accuracy trends of FedGloSS compared to state-of-the-art methods for heterogeneous FL on Cifar10 and Cifar100 respectively. For a clearer understanding, we distinguish between Sam-based and Sgd-based methods depending on the used local optimizer. For a fair comparison, we report FedSmoo with and without the scheduling of $\rho$ (+wp in the figure). For additional details on the scheduling, refer to App. C. FedGloSS consistently achieves the best performances and convergence speedup in each group. In addition, we remind that FedGloSS communicates half the number of bits at each round w.r.t. FedSmoo. Fig. 17 reports the results obtained with ResNet18 on both datasets: FedGloSS consistently achieves the best speed of convergence and final accuracy, with both Sam and Sgd as ClientOpt.
@@ B.4 Effectiveness in Multiple Scenarios and with Several Model Architectures
As already shown in Sec. 6, FedGloSS outperforms the state of the art in terms of speed of convergence, final performance in accuracy, flatness of the solution and communication efficiency. This remains true across multiple datasets (Cifar100, Cifar10, Landmarks-User-160k) and model architectures (CNN, ResNet18, MobileNetv2). The main paper focuses on results in more challenging heterogeneous FL scenarios. The next paragraphs discuss the behavior of FedGloSS in increasingly more homogeneous settings.
FedGloSS in moderately heterogeneous settings.
The choice of the Dirichlet’s concentration parameter $\alpha$ in the main text aligns with the FL literature [1, 4, 20, 53, 54], where $\alpha<1$ is commonly used to evaluate challenging settings. Moreover, for $\alpha>0.6$, differences between methods become less pronounced [61]. To further analyze the impact of varying heterogeneity levels on FedGloSS, we consider $\alpha\in\{1,5,10\}$ on Cifar10. Tab. 10 presents a comparison of FedGloSS against several state-of-the-art FL algorithms. Consistently, FedGloSS achieves the highest accuracy across all tested scenarios, except for $\alpha=5$ with SGD as the local optimizer, where it is outperformed by Scaffold. Notably, even in this case, FedGloSS exhibits the lowest maximum Hessian eigenvalue, highlighting its stability.
Table 10: FedGloSS against SOTA algorithms in increasingly more homogeneous settings on Cifar10, compared in terms of accuracy (%) and maximum Hessian eigenvalue $\lambda_{1}$. Best results in bold. Model: CNN.
Method $\alpha=1$ $\alpha=5$ $\alpha=10$
Accuracy $\lambda_{1}$ Accuracy $\lambda_{1}$ Accuracy $\lambda_{1}$
Sgd-based FedAvg $76.4\pm 0.2$ $68.2\pm 2.7$ $83.7\pm 0.4$ $63.3\pm 2.9$ $82.4\pm 0.1$ $70.6\pm 0.3$
FedProx $76.3\pm 0.2$ $69.5\pm 0.9$ $83.8\pm 0.4$ $63.0\pm 3.6$ $82.4\pm 0.2$ $71.8\pm 2.0$
FedDyn $77.1\pm 0.1$ $64.5\pm 4.9$ $82.3\pm 0.3$ $48.3\pm 1.6$ $82.2\pm 0.1$ $58.1\pm 3.5$
Scaffold $79.6\pm 0.1$ $62.7\pm 2.7$ $84.5\pm 0.2$ $47.4\pm 3.0$ $83.8\pm 0.2$ $58.4\pm 2.9$
[…]
Sam-based FedSam $77.7\pm 0.2$ $25.6\pm 0.02$ $83.5\pm 0.1$ $19.7\pm 0.1$ $83.2\pm 0.2$ $22.1\pm 1.4$
FedDyn $80.5\pm 0.2$ $21.0\pm 0.8$ $82.9\pm 0.4$ $16.3\pm 0.5$ $84.1\pm 0.1$ $17.3\pm 0.3$
FedSpeed $76.2\pm 0.5$ $8.3\pm 0.1$ $83.6\pm 0.1$ $20.3\pm 0.5$ $78.2\pm 0.2$ $8.6\pm 0.2$
FedGamma $70.3\pm 1.0$ $4.8\pm 0.2$ $83.5\pm 0.1$ $16.0\pm 0.6$ $76.4\pm 0.3$ $6.4\pm 0.1$
FedSmoo $85.2\pm 0.2$ $9.8\pm 0.2$ $86.2\pm 0.2$ $6.0\pm 0.2$ $86.8\pm 0.3$ $6.2\pm 0.2$
[…]
@@
Tab. 11 evaluates FedGloSS against the main methods FedAvg, FedSam, FedDyn and FedSmoo in homogeneous FL settings. Here, client gradients are naturally more aligned due to reduced client drift [26]. We thus test FedGloSS with and without ADMM for global consistency. As expected, FedGloSS achieves similar accuracy with or without ADMM, particularly when using SAM as the local optimizer. However, ADMM facilitates convergence to flatter minima (evidenced by lower $\lambda_{1}$ values) by aligning local and global convergence points. Notably, FedGloSS achieves the flattest minima (lowest $\lambda_{1}$) across both datasets, and the best accuracy on the more complex Cifar100. While FedSmoo achieves slightly higher accuracy on Cifar10, FedGloSS finds a flatter minimum and achieves competitive accuracy with significantly lower communication costs (halved).
@@ B.5 Reducing Communication Costs
Tab. 12 extends the analysis presented in Sec. 6.3.4 by evaluating the communication costs across all scenarios considered in this work. Notably, since FedGloSS maintains the per-round communication cost of FedAvg, it remains advantageous even when matching the convergence speed of the best-performing algorithm (FedSmoo), due to its halved communication costs.
@@
Fig. 18 deepens the comparison between FedGloSS and its baseline NaiveFedGloSS, discussed in Sec. 6.4. In particular, this plot reports the accuracy trends of the two methods, showing that NaiveFedGloSS is slightly faster (1.1$\approx$ 1.1$\times$ 1.1 ×) than FedGloSS in Cifar100, while FedGloSS surpasses the speed of the baseline after $\approx 25\%$ of training in Cifar10. However, both methods reach the same accuracy at the of training. In addition, it is to be noted that FedGloSS halves the communication cost w.r.t. NaiveFedGloSS by transmitting half the number of bits at each round, as it eliminates the need to invoke the clients twice to compute the updates. Lastly, our sharpness approximation does not steer the optimization path: models trained with FedGloSS and NaiveFedGloSS end up in the same basin (no loss barrier), with similar flatness (Fig. 19). These insights further support our choice of the efficient strategy of FedGloSS over NaiveFedGloSS.
@@
Figure 19: Loss barriers resulting from interpolating NaiveFedGloSS and FedGloSS’ models, which are found in the same basin. Cifar datasets, CNN. Details on the computation in Sec. D.2.
@@ Appendix C Implementation Details
This section delves into a comprehensive description of the datasets and models utilized throughout this paper, specifying the Deep Learning framework and the hardware employed (Sec. C.1). Additionally, we present the area of the hyper-parameters’ space explored during the fine-tuning process in order to yield optimal results (Sec. C.2).
@@ C.1 Datasets and Models
Tab. 13 provides a comprehensive overview of each dataset’s general statistics. This includes the number of training clients participating in the process and the total number of samples used to construct both the training and test sets.
@@ C.1.1 Cifar10 and Cifar100
We adapted these two well-known image classification datasets to the FL scenario by replicating the splits among clients proposed by [19]. Both datasets are split evenly among 100 clients, thus each of them has access to 500 data samples. This partitioning is performed according to a Latent Dirichlet Allocation (LDA) on the labels. In practice, each local dataset follows a multinomial distribution drawn according to a symmetric Dirichlet distribution with concentration parameter $\alpha$. The higher the value of this parameter is, the more the local datasets resemble a homogeneous scenario, in the limit case $\alpha=0$ each client has access to one only class of images. In our experiments we tested $\alpha\in\{0,0.05,1,5,10,100\}$ and $\alpha\in\{0,0.5,1000\}$ for Cifar10 and Cifar100, respectively. Our choice of $\alpha$ aligns with FL literature [1, 4, 20, 53, 54], where $\alpha<1$ is standard for testing challenging settings. Also, for $\alpha>0.6$, method differences are less apparent [61]. Figs. 21(a) and 21(b) show how data is distributed across clients in all the experimental settings for these two datasets. Both datasets are pre-processed by applying random crops and random horizontal flips.
@@
We trained a Convolutional Neural Network (CNN) inspired by the LeNet-5 architecture [30], as proposed by [20]. The network comprises two 64-channels convolutional layers, both using $5\times 5$ kernels and followed by $2\times 2$ max-pooling layers. This is succeeded by two fully-connected layers with 384 and 192 units, respectively. The final output layer is adapted to the specific number of classes in the dataset.
To explore deeper and more expressive architectures, we also trained a ResNet18 [16] on Cifar100 with $\alpha=0.5$ and Cifar10 with $\alpha=0.05$. We replaced the standard Batch Normalization layers [21] with Group Normalization layers [59] due to their demonstrated effectiveness in handling skewed data distributions in FL [18]. We carried out the experiments with the CNN model using PyTorch [43] and the ResNet18 ones using FedJAX [48].
@@ C.1.2 Landmarks-User-160k
To achieve a comprehensive understanding of the efficacy of the proposed method, a thorough analysis was undertaken utilizing large-scale real-world datasets. Specifically, we use the Landmarks-User-160k dataset [20], a repository encompassing 164,172 images depicting 2028 distinct landmarks, distributed among 1262 clients.
@@
The model employed for training is a MobileNetV2 [50, 20], replacing batch normalization with group normalization layers and pre-trained on ImageNet [8]. The tests on this dataset were carried out on a cluster of NVIDIA A100 40GB, using our FedJAX codebase.
@@ C.2 Hyperparameters
Grid search.
In Tab. 14 we report the training hyperparameters associated to each dataset and model pairing, while Tab. 15 summarizes the hyper-parameters search grid tested for each method (in bold the chosen ones). All runs are averaged over 3 seeds. In addition, following previous works [1, 4, 5], we report the averaged accuracy of the last 100 rounds to reduce the noise typical of heterogeneous FL settings.
Local hyperparameters.
As for local hyperparameters, we tested FedGloSS’ robustness to $E=2$ and $\eta_{l}=0.001$ on Cifar10 with $\alpha=0$. A smaller $\eta_{l}$ had the greatest impact, with FedAvg and FedSam losing 45% and 77% of their performance (more local steps degrade performance in heterogeneous settings), respectively. Notably, FedGloSS loses only 29%. We attribute this to the pseudo-gradient consistently pointing toward higher-loss regions, ensuring a reliable ascent step approximation regardless of local hyperparameters.
Scheduling of $\rho$.
@@ Appendix D Flatness Analysis
D.1 Visualizing 3D Loss Landscapes
We leverage techniques from [31] to visualize the loss landscapes of our models. We adapt their code to work with our specific datasets and network architectures. The process involves calculating the loss function along random directions in the parameter space. This is achieved by perturbing the model’s parameters within a defined range. In our visualizations, we constrain these perturbations to occur within the range of $[-1,1]$ for both the $x$ and $y$ axes. To ensure consistent comparisons across models (e.g., as seen in Fig. 6), we utilize the same set of random directions for all models.
D.2 Visualizing 1D Loss Landscapes
Fig. 19 shows the interpolation of FedGloSS and NaiveFedGloSS’s models. Given their respective weights $\operatorname{\boldsymbol{w}}_{\textsc{FedGloSS}}$ and $\operatorname{\boldsymbol{w}}_{\textsc{NaiveFedGloSS}}$, the interpolation is computed using a coefficient $\gamma$ as
@@ D.3 Hessian Eigenvalues for Flatness Measure
Following prior work [12, 14, 31, 4], we use the spectrum of the Hessian matrix to quantify the flatness of the loss landscape. Here, lower maximum eigenvalues correspond to flatter landscapes, implying less sharpness. To compute these eigenvalues (denoted by $\lambda_{1}$ in the main paper), we employ the stochastic power iteration method [60] with a maximum of 20 iterations, referring to the code of Golmant et al. [15].